# Two Optimizers Are Better Than One: LLM Catalyst Empowers Gradient-Based Optimization for Prompt Tuning

### v3

@@ Abstract
Learning a skill generally relies on both practical experience by doer and insightful high-level guidance by instructor. Will this strategy also work well for solving complex non-convex optimization problems? Here, a common gradient-based optimizer acts like a disciplined doer, making locally optimal update at each step. Recent methods utilize large language models (LLMs) to optimize solutions for concrete problems by inferring from natural language instructions, akin to a high-level instructor. In this paper, we show that these two optimizers are complementary to each other, suggesting a collaborative optimization approach. The gradient-based optimizer and LLM-based optimizer are combined in an interleaved manner. We instruct LLMs using task descriptions and timely optimization trajectories recorded during gradient-based optimization. Inferred results from LLMs are used as restarting points for the next stage of gradient optimization. By leveraging both the locally rigorous gradient-based optimizer and the high-level deductive LLM-based optimizer, our combined optimization method consistently yields improvements over competitive baseline prompt tuning methods. Our results demonstrate the synergistic effect of conventional gradient-based optimization and the inference ability of LLMs. The code is released at https://github.com/guozix/LLM-catalyst.
@@ 1 Introduction
Recent developments of Large Language Models (LLMs) have demonstrated an unprecedented ability to comprehend and generate human-like text, leading to significant breakthroughs in natural language processing Touvron et al. (2023a); Chowdhery et al. (2022). This has led to their adoption in various advanced open-ended applications, where they are being instructed to participate in dialogue (OpenAI et al., 2024), formulate and execute plans (Gupta and Kembhavi, 2022; Gao et al., 2023), solve mathematical problems (Yang et al., 2023), etc., rather than just performing rigid natural language understanding tasks (Wang et al., 2019) like previous language models (Radford et al., 2019; Devlin et al., 2018). The robust capability of LLMs in natural language comprehension and the generation of more nuanced and contextually relevant text provides a foundation for solving complex optimization problems using LLMs.
@@
Humans acquire skills through practical experience and external guidance from instructors. Similarly, solving optimization problems relies on well-designed algorithms incorporating prior knowledge, as well as meticulous procedural implementation. We regard common gradient-based optimizer as disciplined doer. These optimizers are effective in navigating the parameter space through precise, incremental adjustments based on gradient information. However, their local perspective often limits their ability to escape local optima and discover more optimal solutions. LLMs can be used as an high-level instructor, being capable of providing useful guidance for optimization, based on the abundant internal knowledge acquired through massive pre-training, Recent studies have proposed to utilize LLMs as strategy planner or optimizer on concrete optimization tasks. For example, Eureka (Ma et al., 2024) trains agents by reinforcement reward function designed by GPT-4, which can learn various complex skills such as dexterous pen spinning. It shows that LLM can guide the trend of optimized policy on a delicate level.
@@
Employing LLMs for optimization offers unique advantages. The optimization is conducted with natural language interactions, which contributes to two charming properties. First, the implementation of the optimization is code-free. All the steps in the process use dialog-like interactions with LLMs. Secondly, LLMs generate instruction-related outputs by assembling task-related semantic tokens which is hardly constrained by the local optima issue, which is probably encountered by gradient-based optimization. On the other hand, LLM returned solutions, generated by simply query and response interactions, may not be as precise as the results optimized by rigorous step-by-step gradient descent, especially under limited API calling budgets. The local carefulness of gradient-based optimizer and diverse semantic exploration of LLM-based optimizer are complementary to each other, suggesting a collaborative optimization approach.
@@
With this motivation, we propose a combined optimization method that adopts LLMs as an optimizer (MaaO) as a catalyst for the conventional gradient-based optimization. Our proposed optimization approach leverages both the locally rigorous gradient-based optimizer in parameter space and high-level deductive LLM-based optimizer in unconstrained vocabulary space for better performance. To achieve a collaborative training based on the two optimizers, we interleave the conventional training process of gradient-based optimization with interactions with LLM. We first optimize the parameters for only dozens of iterations by gradient optimizer. Then the optimized parameters in the intermediate step, along with their loss and accuracy on the training set, are provided as history trajectory clues for LLM to infer new candidates that are potentially more effective, and, after grabbing the response from LLM, we use the generated results as restarting points of the parameters for subsequent gradient-based optimization iterations. The two optimizers are operated alternately to optimize the parameter collaboratively. The final optimized results are obtained by the gradient optimizer with a stable convergence. Our proposed optimizing strategy only injects several API calls to LLM to the conventional gradient-based training workflow. We instantiate the target problem as prompt optimization and validate that our proposed combined optimizing method leads to consistent improvements.
@@
  • We analyze the limitations of gradient-based optimization, especially the entrapment in local optima, and attribute the issues to the limitation of gradient-based optimizer in the short-sighted local perspective of the parameter space.
@@
  • We test the effectiveness of the proposed combined optimization method on prompt tuning tasks, and it achieves consistent improvements over existing prompt tuning methods, validating the complementary effect of LLM-based and gradient-based optimizers.
@@ 2.1 LLMs and Optimization Problems
LLM’s superior capabilities in grasping intricate linguistic structures and accurate language generation enable it to handle complex optimization problems. Existing works have validated the effectiveness of LLM for solving small-scale mathematical optimization problems (Yang et al., 2023), optimizing prompts (Zhou et al., 2023; Pryzant et al., 2023; Guo et al., 2024; Liu et al., 2023a; Fernando et al., 2023; Diao et al., 2023), searching for network architectures (Chen et al., 2023; Zheng et al., 2023), hyperparameter optimization (Chen et al., 2022) and discovering physical equations (Du et al., 2024).
@@
For prompt optimization, APE (Zhou et al., 2023) proposes a general framework that searches for prompts in the vocabulary space by instructing large language models with task information to obtain better solutions iteratively from the responses of LLMs by analyzing previously found candidates. Based on this framework, APO (Pryzant et al., 2023) proposes that editing prompts by LLM is analogous to conducting gradient descent in the natural language domain. They imitate the gradient-based learning by providing the failure cases to LLM for a semantic "gradient" and updating the prompt in an opposite semantic direction. EVOPROMPT (Guo et al., 2024) also connects LLM-based optimization to traditional algorithms for better explainability. They integrate LLM into the workflow of evolutionary algorithms by instructing LLM to act like evolutionary operators to generate new candidate prompts. The insight that LLM naturally enables an intelligent variation operator is also revealed in LMC (Meyerson et al., 2024) and ELM (Lehman et al., 2022) on image and code generation tasks. BBO(Liu et al., 2023a) searches prompts for vision-language model by merely conversing with LLM following designed strategies and achieves comparable results to white-box gradient-based prompt tuning. However, the performance superiority only holds in one-shot setting. LLM can not effectively optimize to gain more improvements based on more training data.
@@
The results achieved in these approaches demonstrate that LLMs have the potential to be applied as a general-purpose optimizer to more tasks. Some of the works explored the connection between LLM-based inference and existing optimization algorithms (Pryzant et al., 2023; Guo et al., 2024). However, the proposed optimization workflow is still largely based on the inherent ability of LLM. It is still hard for LLM to make decisions as precise as deterministic optimization algorithms, e.g., gradient-based optimizer. And the API calling budget bounded by the high cost of operating super large-scale models also limits the performance of LLM-based optimization.
@@ 2.2 Prompt Tuning for Pre-trained Models
Prompt tuning has emerged as a standard approach for the parameter-efficient adaptation of pre-trained models, aimed at improving their performance in various natural language processing Lester et al. (2021); Li and Liang (2021) and vision-language Zhou et al. (2022b, a); Yao et al. (2024) tasks. Prompt-based tuning of pre-trained models appends learnable embeddings to the original sequence of the data for the input layer or intermediate layer. Fine-tuning the lightweight parameters in the prompt yields comparable performance even to full parameter fine-tuning and transferability (Vu et al., 2022; Su et al., 2022) on various tasks (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2022). Despite its widespread adoption, the conventional prompt tuning technique encounters challenges related to slow convergence and suboptimal optimization Ding et al. (2022), which undermines the effectiveness of prompt tuning in a wider and larger scale of pre-trained models and downstream tasks. We attribute these issues to the complexity of the input embedding space of the pre-trained model, making it challenging to optimize the prompt effectively based on back-propagated gradients in this space.
@@ 3 Method
In this section, we introduce our proposed combined optimization approach that leverages both the local carefulness of gradient-based optimizer and the flexible semantics exploration of LLM-based optimizer. The overview of our method is shown in Figure 1. To elaborate on our proposed method, we instantiate the problem in a prompt tuning scenario. We will describe the general formulation of prompt tuning/optimization and the way of gradient-based prompt tuning in Section 3.1. Next, we analyze the issues that occur in the conventional gradient-based prompt tuning process in Section 3.2, and attribute the problem to the characteristics of gradient-based optimizer that is limited to the local view of the parameter space. Finally we introduce our proposed optimization method that adopts LLM as a catalyst for gradient-based prompt tuning in Section 3.3.
@@ 3.1 General Formulation of Prompt Tuning
In this part, we instantiate the task by prompt tuning for discriminative tasks, i.e., classification. In a general situation, we consider a pre-trained multi-modal model $\mathcal{E}$. Recognizing the classes of input images $I$ or texts $T$ can be achieved by classifying the representations $\mathcal{E}\left(I,T\right)$, encoded by the pre-trained model. We denote the task-specific classifier as $\mathcal{F}\left(\cdot\right)$. The prediction can be obtained by $p\left(\hat{y}|I,T\right)=\mathcal{F}\left(\mathcal{E}\left(I,T\right)\right)$.
@@ 3.2 Analysis on Issues of Gradient-Based Prompt Tuning
Although prompt tuning has become one of the most widely adopted parameter-efficient fine-tuning methods for the adaptation of pre-trained models. The optimization of the prompt still encounters challenges. The prompts converges much slower than other parameters efficient fine-tuning methods, e.g., adapter tuning or even full parameter fine-tuning (Ding et al., 2022), based on the estimated gradients back-propagated through the entire pre-trained model. Another main issue of prompt tuning is that the effectiveness of the learned prompt is sensitive to its initialization values, suggesting that the optimization of the prompt may easily entrapped in local optima due to the complexity of the embedding space of the pre-trained model. Unfortunately, it is challenging to carefully craft initial prompts for every downstream task. To address this issue, Gu et al. (2022) propose to seek a satisfying initialization point for the prompt. However, their method needs to inject soft prompts into the pre-training stage, which limits its application to scenarios where pre-training resource is limited.
@@
To demonstrate the issues more specifically, we analyze some empirical results of gradient-based prompt optimization performed on the one-shot training set of EuroSAT (Helber et al., 2019). We fix the training set for all experiments to eliminate the variance caused by data sampling. We run CoOp (Zhou et al., 2022b) under three random initializations and show the results as indicated by "Random Initialization" in Figure 2. It can be seen that even if we fix the training samples, different random initialization value of the prompt can still bring considerable standard deviation in the results of final learned prompts, indicating a large performance gap (up to 9 percent of accuracy) between different seeds. If we manually initialize the prompt as a prompt template "a photo of a", which is used in Radford et al. (2021), the final variance gets smaller but the absolute performance shows a slight decline. Prior knowledge contained in manual prompts brings merits, providing better results at the starting phase of the training, but lacks proper flexibility for enhancement of final learned prompts. Our method adds marginal steps of optimization based on the collaboration of gradient-based optimizer and MaaO at the start of the training workflow, which results in both lower standard deviation and better absolute performance.
@@
The high sensitivity of prompt tuning results according to different initialization values indicate the complexity of the input embedding space, where gradient-based optimizer only leads to suboptimal converged parameters based on gradient information in a short-sighted local perspective, hardly taking the semantics of the prompt and the overall task information into account. To mitigate the limitations of gradient-based optimizer, we leverage LLM as an unconstrained vocabulary space prompt optimizer based on textual semantic information of the task and previously found prompts.
@@
16:     Reinitialize: $\widetilde{\bm{p}}_{\bm{\theta}}\leftarrow\mathcal{V}(\widetilde{p})$, $\mathcal{P}\leftarrow\varnothing$
@@
3.3 Catalyzing Prompt Tuning by Using LLM as a Prompt Optimizer

### v4

# LLM as a Complementary Optimizer to Gradient Descent: A Case Study in Prompt Tuning

@@ Abstract
Mastering a skill generally relies on both hands-on experience from doers and insightful, high-level guidance by mentors. Will this strategy also work well for solving complex non-convex optimization problems? Here, a common gradient-based optimizer acts like a disciplined doer, making locally optimal updates at each step. Large Language Models (LLMs) can also search for better solutions by inferring from natural language instructions, akin to a high-level mentor. In this paper, we show that these two participators are complementary to each other and can effectively collaborate as a combined optimization framework. The collaborative optimization is achieved by alternating between the gradient-based and LLM-based optimizers. We instruct LLMs to generate possibly improved solutions by taking parameter trajectories recorded during the previous stage of gradient-based optimization into account. Inferred results of LLMs are used as restarting points for the next stage of gradient optimization. We verify the effectiveness of this optimization framework on prompt tuning. By leveraging both the locally rigorous gradient-based optimizer and the high-level deductive LLM-based optimizer, the combined optimization method consistently yields improvements over competitive baselines on a variety of tasks. Our results demonstrate the synergistic effect of conventional gradient-based optimization and the inference ability of LLMs.
The code will be made publicly available.
@@ 1 Introduction
Humans acquire skills through practical experience and external guidance from mentors. Similarly, solving optimization problems relies on well-designed algorithms incorporating prior knowledge, as well as meticulous procedural implementation. Practically, gradient-based algorithms have almost become the default choice for solving optimization problems in various machine learning models. We regard the gradient-based optimizers as disciplined doers that are effective in navigating the parameter space through precise, incremental adjustments based on gradient information. However, their local perspective often limits their ability to escape local optima and discover more optimal solutions.
In this work, we proposed an optimization method using LLMs as optimization instructors to provide high-level guidance for gradient-based optimizers. The basis of LLMs capable of solving optimization problems lies in their ability to comprehend and generate nuanced and contextually relevant text. Recent studies have proposed to utilize LLMs as strategy planners or optimizers in concrete optimization tasks. For example, Eureka (Ma et al., 2024) trains agents by reinforcement reward function designed by GPT-4, which can learn complex skills such as dexterous pen spinning. It shows that LLM can guide the trend of optimized policy on a delicate level.
@@
Employing LLMs for optimization offers unique advantages. The optimization is conducted with natural language interactions, which contributes to two charming properties. First, the implementation of the optimization is code-free. The optimization process only involves natural language instruction-response interactions with LLMs. Secondly, LLMs generate instruction-related outputs by assembling task-related semantic tokens that are difficult to discover through continuous gradient-based learning. The generation results can be diverse and hardly limited by the local optima issue, which is often encountered by gradient-based optimization. The solutions discovered by LLM with a lower loss value have more possibilities to optimize to a better convergence point.
On the other hand, LLM-based optimization faces instability issues since it has no lexical constraints. LLMs analyze the problem on the semantic level and response in vocabulary space. The generated solutions may not be as precise as the results optimized by rigorous step-by-step gradient descent in the parameter space, especially under limited LLM API calling budgets. Existing LLM-based methods, such as  Liu et al. (2023a), need to set multiple search trials to find a promising result. We instruct LLMs with the optimization trajectory of gradient-based optimizers, which is converted to natural language format, grounding LLMs to a more promising sub-region of the vocabulary space. The local carefulness of gradient-based optimizer and diverse semantic exploration of LLM-based optimizer are complementary to each other, suggesting a collaborative optimization approach.
@@
With this motivation, we propose an optimization method that combines the conventional gradient-based optimizer and LLM optimizer. The optimization approach leverages both the locally rigorous gradient-based optimizer in parameter space and high-level deductive LLM-based optimizer in unconstrained vocabulary space for better performance. To achieve collaborative training based on the two optimizers, we interleave the conventional training process of gradient-based optimization with interactions with LLM. First, we optimize the parameters for only dozens of iterations using a gradient optimizer. Then the optimized parameters in the intermediate step, along with their loss and accuracy on the training set, are provided as history trajectory clues for LLM to infer new candidates that are potentially more effective. After grabbing the response from LLM, we use the generated results as restarting points of the parameters for subsequent gradient-based optimization iterations. The two optimizers are operated alternately to optimize the parameter collaboratively. The final optimized results are obtained by the gradient optimizer with a stable convergence. Our proposed optimizing strategy only injects several API calls to LLM to the conventional gradient-based training workflow.
We study the effectiveness of the combined optimization in a widely considered prompt tuning framework. Specifically, we optimize textual prompts for language and vision-language pre-trained models. Tuning such prompts is shown to bring significant performance improvements for the adaptation of pre-trained models (Zhou et al., 2022b; Lester et al., 2021). However, optimizing the prompt in the input discrete vocabulary space or word embedding space is not an easy problem for conventional gradient-based optimizers that is widely adopted. We validate that the proposed collaborative optimization framework leads to consistent improvements in prompt optimization across a variety of tasks and optimizer LLMs.
@@
  • Based on the textual prompt optimization problem, we showcase the limitations of gradient-based optimization, especially the entrapment in local optima, and attribute the issues to the limitation of gradient-based optimizer in the short-sighted local perspective of the parameter space.
@@
  • We test the effectiveness of the proposed combined optimization method on prompt tuning tasks, and it achieves consistent improvements over existing competitive baseline methods, validating the complementary effect of LLM-based and gradient-based optimizers.
@@ 2.1 LLMs and Optimization Problems
Recent developments of Large Language Models (LLMs) have demonstrated an unprecedented ability to comprehend and generate human-like text, leading to significant breakthroughs in natural language processing (Touvron et al., 2023a; Chowdhery et al., 2022). The robust capability of LLMs in natural language comprehension and the generation of more nuanced and contextually relevant text provides a foundation for various advanced open-ended applications, where they are being instructed to participate in dialogue (OpenAI et al., 2024), formulate and execute plans (Gupta & Kembhavi, 2022; Gao et al., 2023), writing codes (Ma et al., 2024), etc. LLMs’ rich prior knowledge and reasoning ability open the way to addressing practical optimization problems for real-world applications. Existing works have validated the effectiveness of LLM for solving small-scale mathematical optimization problems (Yang et al., 2023), optimizing prompts (Zhou et al., 2023; Pryzant et al., 2023; Guo et al., 2024; Liu et al., 2023a; Fernando et al., 2023; Diao et al., 2023; Zhang et al., 2023), searching for network architectures (Chen et al., 2023; Zheng et al., 2023), hyperparameter optimization (Chen et al., 2022) and discovering physical equations (Du et al., 2024).
@@
In terms of prompt optimization, APE (Zhou et al., 2023) proposes to use LLMs to generate and select natural language prompts by instructing LLMs with task definitions and targets. LLMs can obtain better solutions iteratively by analyzing previously found candidates. APO (Pryzant et al., 2023) proposes that editing prompts by LLM is analogous to conducting gradient descent in the natural language domain. They imitate the gradient-based learning by providing the failure cases to LLM for a semantic "gradient" and updating the prompt in an opposite semantic direction. EVOPROMPT (Guo et al., 2024) also connects LLM-based optimization to traditional algorithms for better explainability. They integrate LLM into the workflow of evolutionary algorithms by instructing LLM to act like evolutionary operators to generate new candidate prompts. The insight that LLM naturally enables an intelligent variation operator is also revealed in LMC (Meyerson et al., 2024) and ELM (Lehman et al., 2022) on image and code generation tasks. Liu et al. (2023a) searches prompts for the vision-language model by conversing with LLM following designed strategies and achieves comparable results to white-box gradient-based prompt tuning.
@@
The results achieved in these approaches demonstrate that LLMs can be applied as a general-purpose optimizer for optimization tasks. Although some of them (Pryzant et al., 2023; Guo et al., 2024) explored the connection between LLM-based inference and conventional optimization algorithms, e.g., gradient descent, evolutionary computing. However, the proposed optimization workflows are still largely based on the inherent ability of LLM, which leads to inadequate data utilization and suboptimal performance. For example, in Liu et al. (2023a), the performance superiority only holds in the one-shot training set and LLMs can not effectively optimize to gain more improvements based on more training data. It shows that the interaction format of natural language makes it hard for LLMs to optimize as precisely as numerical optimization algorithms, e.g., gradient-based optimizers. Besides, the API calling budget bounded by the high cost of operating super large-scale models also limits the performance of LLM-based optimization. This motivates us to design a collaborative optimization method to achieve better optimization performance by combining both the results of LLMs’ high-level reasoning and the stable convergence of conventional gradient-based optimizers.
@@ 2.2 Prompt Tuning for Pre-trained Models
Prompt tuning has emerged as a standard approach for the parameter-efficient adaptation of pre-trained models, aimed at improving their performance in various natural language processing (Lester et al., 2021; Li & Liang, 2021) and vision-language (Zhou et al., 2022b; a; Yao et al., 2024) tasks. Prompt-based tuning of pre-trained models appends learnable embeddings to the original sequence of the data for the input layer or intermediate layer. Fine-tuning the lightweight parameters in the prompt yields comparable performance even to full parameter fine-tuning and transferability (Vu et al., 2022; Su et al., 2022) on various tasks (Lester et al., 2021; Li & Liang, 2021; Liu et al., 2022). Despite its widespread adoption, the conventional prompt tuning technique encounters challenges related to slow convergence and suboptimal optimization (Ding et al., 2022), which undermines the effectiveness of prompt tuning in a wider and larger scale of pre-trained models and downstream tasks. We attribute these issues to the complexity of the input embedding space of the pre-trained model, making it challenging to optimize the prompt effectively based on back-propagated gradients in this space.
@@ 3 Method
In this section, we introduce our proposed combined optimization approach that leverages both the local carefulness of gradient-based optimizer and the flexible semantics exploration of LLM-based optimizer. The overview of our method is shown in Figure 1. We instantiate the problem in a prompt tuning scenario to elaborate on our proposed method. We will describe the general formulation of prompt tuning/optimization and the way of gradient-based prompt tuning in Section 3.1. Next, we analyze the issues that occur in the conventional gradient-based prompt tuning process in Section 3.2, and attribute the problem to the characteristics of gradient-based optimizer that is limited to the local view of the parameter space. Finally, we introduce our proposed combined optimization method in Section 3.3.
@@ 3.1 General Formulation of Prompt Tuning
In this part, we instantiate the task by prompt tuning for discriminative tasks, i.e., classification. In a general situation, we consider a pre-trained multi-modal model $\mathcal{E}$. The classes of input images $I$ or texts $T$ can be recognized by classifying the representations $\mathcal{E}\left(I,T\right)$, encoded by the pre-trained model. We denote the task-specific classifier as $\mathcal{F}\left(\cdot\right)$. The prediction can be obtained by $p\left(\hat{y}|I,T\right)=\mathcal{F}\left(\mathcal{E}\left(I,T\right)\right)$.
@@ 3.2 Analysis on Issues of Gradient-Based Prompt Tuning
Although prompt tuning has become one of the most widely adopted parameter-efficient fine-tuning methods for the adaptation of pre-trained models. The optimization of the prompt still encounters challenges. The prompts converge much slower than other parameters efficient fine-tuning methods, e.g., adapter tuning or even full parameter fine-tuning (Ding et al., 2022), based on the estimated gradients back-propagated through the entire pre-trained model. Another main issue of prompt tuning is that the effectiveness of the learned prompt is sensitive to its initialization values, suggesting that the optimization of the prompt may easily entrapped in local optima due to the complexity of the embedding space of the pre-trained model. Unfortunately, it is challenging to carefully craft initial prompts for every downstream task. To address this issue, Gu et al. (2022) propose to seek a satisfying initialization point for the prompt. However, their method needs to inject soft prompts into the pre-training stage, which limits its application to scenarios where pre-training resources are limited.
@@
To demonstrate the issues more specifically, we analyze some empirical results of gradient-based prompt optimization performed on the one-shot training set of EuroSAT (Helber et al., 2019). We fix the training set for all experiments to eliminate the variance caused by data sampling. We run CoOp (Zhou et al., 2022b) under three random initializations and show the results as indicated by "Random Initialization" in Figure 2. It can be seen that even if we fix the training samples, different random initialization values of the prompt can still bring considerable standard deviation in the results of final learned prompts, indicating a large performance gap (up to 9 percent of accuracy) between different seeds. If we manually initialize the prompt as a prompt template "a photo of a", which is used in Radford et al. (2021)’s work, the final variance gets smaller but the absolute performance shows a slight decline. Prior knowledge contained in manual prompts brings merits, providing better results at the starting phase of the training, but lacks proper flexibility for enhancement of final learned prompts. Our method adds marginal steps of optimization based on the collaboration of gradient-based optimizer and MaaO at the start of the training workflow, which results in both lower standard deviation and better absolute performance.
@@
The high sensitivity of prompt tuning results according to different initialization values indicates the complexity of the input embedding space, where gradient-based optimizer only leads to suboptimal converged parameters based on gradient information in a short-sighted local perspective, hardly considering the semantics of the prompt and the overall task information. To mitigate the limitations of the gradient-based optimizer, we leverage LLM as an unconstrained vocabulary space prompt optimizer based on textual semantic information of the task and previously found prompts.
@@
16:     Reinitialize: ${\bm{p}}_{\bm{\theta}}\leftarrow\mathcal{V}(\widetilde{p})$, $\mathcal{P}\leftarrow\varnothing$
@@
3.3 A Collaborative Optimization Method by Using LLM as a Prompt Optimizer